OpenStudy (anonymous):

I need some help please! (: 4(b – 3) – b = 7 b + 10

OpenStudy (anonymous):

4(b-3)-b=7b+10 4b-12-b=7b+10 3b-12=7b+10 -3b -3b -12=4b+10 -10 -10 -22=4b /4 /4 B=-5 and a 1/2 (= if I helped reward and medal please..

OpenStudy (anonymous):

Thanks so much! (:

OpenStudy (anonymous):

welcome.